Find the critical value za/2 that corresponds to the given confidence level of 88%

1
Expert's answer
2022-06-01T11:28:36-0400

For confidence "(1-\alpha)(100\%)=88\%, \alpha=0.12."


"z_{\alpha/2}=z_{0.12/2}=\pm 1.5548"